Opened 5 years ago

Last modified 4 years ago

#529 new flight processing

GB12/03, flight day 169/2014, Loch Lomond

Reported by: knpa Owned by: tec
Priority: alpha 4 medium Milestone: 2014 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by tec)

Data location: /users/rsg/arsf/arsf_data/2014/flight_data/uk/GB12_03-2014_169_Loch_Lomond

Data arrived from ARSF via network transfer on 18/06/2014

GB12/03 Scientific objective: Improve lake remote sensing algorithms

GB12/03 Priority: 9

PI: Peter Hunter

Sensors:

Camera (Not Requested), delivered 14/08/2014, Archvied 13/05/2015
Fenix (Requested), delivered 14/08/2014, Archvied 13/05/2015
LiDAR (Not Requested), delivered 14/08/2014 , Archvied 13/05/2015
FW LiDAR (Not Requested), delivered 14/08/2014, Archvied 13/05/2015
OWL (Not Requested)

Change History (37)

comment:1 Changed 5 years ago by tec

Beginning navigation processing.

Last edited 5 years ago by tec (previous) (diff)

comment:2 Changed 5 years ago by tec

KILN Basestation Data
Latitude56 27 17.79077
Longitude-4 19 14.64320
El Height232.139m


LOCG Basestation Data
Latitude56 00 48.73288
Longitude-5 26 45.81319
El Height72.112m

Had to change the elevation mask to 15.5 to get an ambiguity fix of 2 for the majority of the flightlines. Flightlines 1, 12 and the 2 cross flightlines have an ambiguity fix of 1.

Last edited 5 years ago by tec (previous) (diff)

comment:3 Changed 5 years ago by knpa

  • Description modified (diff)

comment:4 Changed 5 years ago by tec

Started hyperspectral processing.

comment:5 Changed 5 years ago by dap

Beginning LiDAR processing.

comment:6 Changed 5 years ago by tec

Hyperspectral SCT Values

Flightline FENIX
1 1.02
2 0.97
3 0.90
4 0.90
5 0.00
6 0.95
7 1.00
8 0.09
9 1.00
10 0.09
11 -0.07
12 0.90
13 0.90
14 0.90
15 -0.20

comment:7 Changed 5 years ago by dap

Low return values from ALSPP, potentially due to large body of water.

comment:8 Changed 5 years ago by tec

Fenix Awaiting delivery check

Last edited 5 years ago by tec (previous) (diff)

comment:9 Changed 5 years ago by dap

Roll error of -0.005672596 seems to be consistent across all LiDAR lines.

comment:10 Changed 5 years ago by tec

Started RCD Processing.

Last edited 5 years ago by tec (previous) (diff)

comment:11 Changed 5 years ago by tec

Removed image 18114722100001G4 due to over exposure

comment:12 Changed 5 years ago by tec

Finished processing RCD. Awaiting delivery check.

comment:13 Changed 5 years ago by dap

Classifying FW lidar data

comment:14 Changed 5 years ago by dap

Finished lidar and full waveform processing. Awaiting delivery check.

Last edited 5 years ago by dap (previous) (diff)

comment:15 Changed 5 years ago by dap

RCD delivery check completed, with no problems found.

comment:16 Changed 5 years ago by knpa

  • Owner set to tec

Beginning Fenix delivery check.

comment:17 Changed 5 years ago by knpa

Fenix delivery check complete.
Following issues need addressing:

  • ReadMe file table is messed up between pages. Easiest fix is to put \newline above the table in the .tex
  • Screenshots directory is empty. Need to put these images in ReadMe too.
  • Logsheet is hand-written version - should create one with logsheetGenerator.py
  • Remove fodis dir as not needed with Fenix
  • Need to unix2dos sensor_FOV_vectors/fenix_viewvector_list.txt
  • 447 is blank. Need to add this to ReadMe. Ask mark1 if should go in standard text.
  • mention underflows over water (most flightlines) in data quality section

comment:18 Changed 5 years ago by tec

Fenix DC issues:

  • Fixed readme, added \newpage to move the table to the next page.
  • Removed fodis dir.
  • Fixed fenix_viewvector_list.txt line terminators.
  • Fixed ReadMe and added data quality remarks.
  • Fixed Logsheet
  • Fixed screenshots and mosaic.

think thats all. Ready to be DC'd again.

comment:19 Changed 5 years ago by tec

Starting LiDAR DC.

comment:20 Changed 5 years ago by tec

Missing PDF: doc/full_waveform.pdf
Missing Directory: fw_extractions

Missing CRLF Line Terminators: flightlines/discrete_laser/ascii/*
running: find -name '*.txt' -print0 | xargs -0 unix2dos
should do it, not sure if ascii files should have CRLF

LiDAR false positive classification on bushes (pretty much most of the bushes), needs to be fixed.

Removed GB12_03-2014_169-intensity.tif
Removed logsheet.tex
Removed logsheet.xml
Renamed Read_Me-20140730_01.pdf -> Read_Me-20140730.pdf

Finished LiDAR DC

Last edited 5 years ago by tec (previous) (diff)

comment:21 Changed 5 years ago by dap

Fixing false noise.

comment:22 Changed 5 years ago by dap

Fixed problems with LiDAR data:

  • Missing pdf has been copied to the delivery directory
  • ascii files changed to CRLF line terminators
  • Reclassified LAS files

Ready for re-delivery check

comment:23 Changed 5 years ago by tec

Cool, LiDAR DC Done, no issues. Classification looks good.

comment:24 Changed 5 years ago by dap

Found that vectors weren't included in the delivery data. Have contacted the PI regarding this, awaiting reply.

comment:25 Changed 5 years ago by dap

Hyperspectral, LiDAR (with full waveform) and camera data dispatched to Peter Hunter.

comment:26 Changed 5 years ago by dap

  • Description modified (diff)

comment:27 Changed 5 years ago by dap

  • Component changed from Processing: general to Archiving

comment:28 Changed 5 years ago by tec

Archiving
Starting archiving

comment:29 Changed 5 years ago by tec

  • Description modified (diff)

Archiving
Done confirmed as of 13/05/2015 12:10

comment:30 Changed 4 years ago by mark1

fenix

Fenix data needs reprocessing due to fault with sensor. Incorrect SWIR times recorded in raw files. Have changed tint2 value in hdr files from 14.4 to 2.15 as recommended by Specim. Data are also binned spectrally.

Data have been reprocessed.

comment:31 Changed 4 years ago by mark1

Fenix

Have updated sct for line 10 to 0.96.

comment:32 Changed 4 years ago by asm

RE-DC for Fenix
Started delivery check.
Reprocessed and re-checked from /data/molucca2/scratch/arsf-temporary/

comment:33 Changed 4 years ago by asm

Fenix delivery check for re-processing completed, everything O.K.

comment:34 Changed 4 years ago by dac

Fenix

Started zipping mapped files on molucca2, will copy back once zipped.

comment:35 Changed 4 years ago by dac

Fenix

Started zipping mapped files on molucca2, will copy back once zipped.

comment:36 Changed 4 years ago by dac

Fenix

Sent reprocessed data on USB hard drive to Peter Hunter on 31/07/2015

comment:37 Changed 4 years ago by dac

Note: See TracTickets for help on using tickets.